You have just purchased a share of stock for

$ 18.86 The company is expected to pay a dividend of

$ 0.58 per share in exactly one year. If you want to earn a

10.6 % return on your​ investment, what price do you need if you expect to sell the share immediately after it pays the​ dividend?
Answer #1

Given Purchased a share of stock P0 =18.86

Expected Dividend D1 =0.58

Return on investment Re. =10.6%

Now you need the price expected to sell the share immediately after pays the dividend P1 =P1

Formula for calculating the price of a share is

P0 = (P1 + D1)/(1 + Re)

18.86 =(P1 +0.58)/(1+10.6%)

P1+0.58 =18.86 * 1.106

P1 =20.85916 - 0.58

P1 = $20.27916 =$20.28

After one year the price is $20.28